OpenStudy (anonymous):

f(x)=(1-x/20)^5, f''(40)=?

OpenStudy (anonymous):

Take the derivative; \[f'(x)= d/dx(1-x/20)^5 \implies f'(x) = 5(-1/20)(1-x/20)^4 \] \[\implies f'(x) = \frac{-1}{4}(1-\frac{x}{20})^4\] Now plug in 40.

OpenStudy (anonymous):

not an option. and it doesn't look right

OpenStudy (anonymous):

I didn't plug in 40...This website isn't for you to cheat on your hw in the first place. Just plug in the number 40 for x...the problem is done for you...

OpenStudy (anonymous):

no im looking for help with taking the derivative, the derivative is wrong because if i plug in the 40 it doesn't give me the answer listed in the options

OpenStudy (anonymous):

ur function is hard to read. is it x/20? or (1-x)/20? clarify. ty

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=derivative+of+(1-x%2F20)%5E5 Tell me its wrong...

OpenStudy (anonymous):

By staring at malevolence19's work, it is correct: There for, ur function is (1-x)/20, not 1-(x/20).

OpenStudy (anonymous):

-1/4

OpenStudy (anonymous):

(always show work) ppl cant learn if u just give answers == ()